In right triangle ABC, the measure of angle A is 45°. If the hypotenuse of the triangle has length 5, what is the length of segment BC?

Answer by ewatrrr(24785)


Isosceles Right Triangle
x^2 + x^2 = 5^2 (Tossing out the negative solution for unit measure)
x^2 = 25/2
x = 5√2/2 = BC